find f(3)/g(1)
f(x)=x^2-x and g(x)=3x+2

Answer by checkley77(12844)

find f(3)/g(1)
f(x)=x^2-x and g(x)=3x+2
f(3)=3^3-3=9-3=6 ans.
g(1)=3*1+2=3+2=5 ans.
f(3)/g(1)=6/5 ans.